OpenStudy (anonymous):

The circumference of a circle is 8π cm. What is the DIAMETER of the circle?

OpenStudy (anonymous):

i used this formula 8=pi*d

OpenStudy (shayaan_mustafa):

Circumference is given as C=(2*pi)(r) where r=radius (Half of diameter) i.e. r=d/2 so circumference becomes. C=(2*pi)(d/2)

OpenStudy (anonymous):

i still dont understand what the diameter would be

OpenStudy (shayaan_mustafa):

 radius is the half of diameter i.e. r=d/2 and diameter is the twice of radius i.e. d=2r